I have a homework question on an oblique asymptote. The question is:
f(x)=x^2+5x-4 over x+2. I got the V.A. x=-2 and the O.A. y= x+3. I got my y=1x+3. My question is the graphing part. I do not know which direction to graph and plot the points. Can you please help me?

Answer by Edwin McCravy(20086)

I have a homework question on an oblique asymptote. The question is:
f(x)=x^2+5x-4 over x+2. I got the V.A. x=-2 and the O.A. y= x+3. I got my y=1x+3. My question is the graphing part. I do not know which direction to graph and plot the points. Can you please help me?



The dark line is the vertical asymptote x=-2 and the green
line is the oblique asymptote y = x+3

Edwin